Greg is 13 years older than his brother. in nine years, Greg will be twice as old as his brother. How old are they now
Answer by Menjax(62)

Let G = Greg's age
Let B = Brother's age
G = B + 13
(The brother is thirteen years younger than Greg, so If you add thirteen years to the brother's age, Greg and his bro are the same age.)
(In 9 years (Greg get's 9 years older) Greg will be twice his brother's age)
G + 9 = 2B
Using simultaneous equations (substitution method)
(B + 13) + 9 = 2B
B + 22 = 2B
22 = B
Sub that into equation 1
G = 22 + 13
G = 35
So B=22 and G=35